You are saving for retirement. On April 1, you deposit $510 in your target-date fund. By May 15, your initial deposit is worth $538 and you deposit an additional $284 at that time. On August 15, your account shows a value of $817 and you make a further deposit of $280. On November 30, your account shows a balance of $1,131. What is your time-weighted return (in percent) for the period April 1 - November 30? Answer to two decimals. Carry intermediate calculations to six decimals.
Answer #1

Answer:

Holding period return from Apr 1 to May 15 = (538 - 500) / 500 = 7.600000%

Holding period return from May 15 to Aug 15 = (817 - (538 + 284)) / (538 + 284) = - 0.608273%

Holding period return from Aug 15 to Nov 30 = (1131 - (817 + 280)) / (817 + 280) = 3.099362%

Time-weighted return (in percent) for the period = ((1 + 7.600000%) * (1 - 0.608273) * (1 + 3.099362)) - 1 = 10.26%

Time-weighted return (in percent) for the period April 1 - November 30 = 10.26%